OpenStudy (anonymous):

Find the value of 7 times x squared all over the square root of x to the fifth power. 7 times the square root of x, all over x. 7 over x. 7 times the square root of x. 7 times x all over the square root of x.

OpenStudy (anonymous):

Alright, so we start with 7x^2/sqrt(x)^5 This can be simplified to 7x^2/x^(5/2) Then, separate the denominator to x^2*x^(1/2) You can then cancel out the x^2 in the numerator and the x^2 in the denominator, leaving you with x^(1/2) which is the same thing as sort(x) So, the final answer should be 7/sqrt(x)

OpenStudy (anonymous):

So which choice would that be?

OpenStudy (anonymous):

NEw best answer for who answers

OpenStudy (anonymous):

Any takers?

OpenStudy (anonymous):

got the answer of 7/sqrt(x). from the choices, what do you think would equal this when simplified?

OpenStudy (anonymous):

just 7 over x?

OpenStudy (anonymous):

or is that the answer?? i got 7 over x i just don't know if that is right

OpenStudy (anonymous):

what happens when you have x^(1/2)/x ?

OpenStudy (anonymous):

It stays the same

OpenStudy (anonymous):

no. you get 1/(x^(1/2)) because (1/2)[from the numerator] - 1[from the denominator] = -1/2 x^(-1/2) = 1/(x^(1/2))

OpenStudy (anonymous):

so what would be the wording for that?

OpenStudy (anonymous):

Oh okayyy so the final answer is 7/sqrt(x)?

OpenStudy (anonymous):

yes. just like what trogdortheburninator said. but i assume we are looking for the answer in sentence form. so that would be the first.

OpenStudy (anonymous):

okay thank you so much!

OpenStudy (anonymous):

I am about to put another one up can you walk me through it?